Solve the following linear equations involving 1 variable

7Y-1=23-5Y

3[2-4(X-1)]=3-4(X+2)

Thanks so much for your help!!!

Answer by stanbon(75887)

7Y-1=23-5Y
-----------
12Y = 24
y = 2
-----------------
3[2-4(X-1)]=3-4(X+2)
3[2-4X+4] = 3-4X-8
18-12X = -5-4X
8X = 23
X = 23/8
=========================
Cheers,
Stan H.